### v1

@@ 3.1 Well-posedness of the Navier–Stokes equations with membrane boundary conditions
where we have used (21a) and (21b).
In the context of the fixed point analysis of (17a)-(17b), for a given $\zeta=\zeta_{0}+\Theta\in Z$ we define the linear functional $G_{\zeta}\in(Q\times W)^{\prime}$ as follows
$\langle G_{\zeta},(q,\xi)\rangle:=\langle\xi,g(\zeta)\rangle_{\Gamma_{\mathrm{m}}}\quad\forall(q,\xi)\in Q\times W,$
(which, thanks to (5) satisfies $\|G_{\zeta}\|_{(Q\times W)^{\prime}}=g_{2}$). Similarly, for a given $\boldsymbol{w}_{0}\in\boldsymbol{\mathrm{H}}_{0}$ we define the linear functional $F_{\boldsymbol{w}_{0},\boldsymbol{U}}\in\boldsymbol{\mathrm{H}}_{0}^{\prime}$:
@@
where $\boldsymbol{U}\in\boldsymbol{\mathrm{H}}^{1}(\Omega)$ is the divergence-free lifting defined in (10). Then, there holds (see [13, Lemma 16])
$\|\boldsymbol{U}\|_{1,\Omega}\lesssim\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}.$ (22)
@@
For known liftings $\boldsymbol{U},\Theta$ and given $\boldsymbol{w}_{0}\in\boldsymbol{\mathrm{H}}_{0}$ and $\zeta_{0}\in Z_{0}$, there exists a unique $(\boldsymbol{u}_{0},p,\lambda)\in\boldsymbol{\mathrm{H}}_{0}\times Q\times W$ such that {linenomath}
$a(\boldsymbol{u}_{0},\boldsymbol{v})+b(\boldsymbol{v},(p,\lambda))$ $=F_{\boldsymbol{w}_{0},\boldsymbol{U}}(\boldsymbol{v})\quad\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{0},$ (23a)
$b(\boldsymbol{u}_{0},(q,\xi))$ $=G_{\zeta}(q,\xi)\quad\forall(q,\xi)\in Q\times W.$ (23b)
@@
$\|\boldsymbol{u}_{0}\|_{1,\Omega}$ $\leq\frac{1}{\underline{\alpha}}\left[\|F_{\boldsymbol{w}_{0},\boldsymbol{U}}\|_{\boldsymbol{\mathrm{H}}_{0}^{\prime}}+\frac{\underline{\alpha}+\|a\|}{\beta}g_{2}\right],$ (24a)
$\|(p,\lambda)\|_{Q\times W}$ $\leq\frac{1}{\beta}\left[\left(1+\frac{\|a\|}{\underline{\alpha}}\right)\|F_{\boldsymbol{w}_{0},\boldsymbol{U}}\|_{\boldsymbol{\mathrm{H}}_{0}^{\prime}}+\frac{\|a\|(\underline{\alpha}+\|a\|)}{\underline{\alpha}\beta}g_{2}\right].$ (24b)
@@
$\|G_{\zeta}\|_{(Q\times W)^{\prime}}=g_{2}.$ (25)
Also, after using triangle inequality together with the boundedness properties in (12a) and (12d), we obtain
$|F_{\boldsymbol{w}_{0},\boldsymbol{U}}(\boldsymbol{v})|$ $=|-\tilde{a}(\boldsymbol{w}_{0}+\boldsymbol{U},\boldsymbol{w}_{0}+\boldsymbol{U},\boldsymbol{v})-a(\boldsymbol{U},\boldsymbol{v})|,$
$\leq C_{i}\rho_{0}\|\boldsymbol{w}_{0}+\boldsymbol{U}\|_{1,\Omega}^{2}\|\boldsymbol{v}\|_{1}+\mu_{0}\|\boldsymbol{U}\|_{1,\Omega}\|\boldsymbol{v}\|_{1,\Omega},$
@@
Let us remark that the unique Stokes velocity from Lemma 3.3 will be in the space $\boldsymbol{\mathrm{X}}^{g}$. We now derive a fixed-point strategy for (17a)–(17b). Let us define the map
@@
$0<4R_{0}<1-\sqrt{1-4(\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}^{2}+g_{2}+2\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}})}.$ (26)
@@
$a(\boldsymbol{u}_{01}-\boldsymbol{u}_{02},\boldsymbol{v})+b(\boldsymbol{v},(p_{1}-p_{2},\lambda_{1}-\lambda_{2}))$ $=F_{\boldsymbol{w}_{1},\boldsymbol{U}}(\boldsymbol{v})-F_{\boldsymbol{w}_{2},\boldsymbol{U}}(\boldsymbol{v})\quad\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{0},$ (27a)
$b(\boldsymbol{u}_{01}-\boldsymbol{u}_{02},(q,\xi))$ $=0\quad\forall(q,\xi)\in Q\times W.$ (27b)
@@
On the other hand, in (3.8) we can use as test functions $\boldsymbol{v}=\boldsymbol{u}_{01}-\boldsymbol{u}_{02}$, $q=p_{1}-p_{2}$, $\xi=\lambda_{1}-\lambda_{2}$ and subtract the two equations to obtain
@@
$L_{\mathcal{J}_{1}}=\frac{1}{\underline{\alpha}^{2}}(2R_{0}+\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}).$ (28)
@@
$L_{\mathcal{J}_{1}}=2R_{0}+\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}<\underline{\alpha}^{2}.$ (29)
@@ 3.2 Well-posedness of the advection–diffusion equation
$\leq\frac{1}{\underline{\alpha}}\left[\|F_{\boldsymbol{w}_{0},\boldsymbol{U}}\|_{\boldsymbol{\mathrm{H}}_{0}^{\prime}}+\frac{\underline{\alpha}+\|a\|}{\beta}g_{2}\right](1+\|\theta_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}})+\|\theta_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}},$ (30)
@@ 3.3 Fixed-point analysis for the coupled flow–transport problem
[…]
where $T:\boldsymbol{\mathrm{H}}\to\boldsymbol{\mathrm{H}}$ is defined as $\boldsymbol{u}\mapsto T(\boldsymbol{u})=({S}_{I}\circ\tilde{S})(\boldsymbol{u}_{0}+\boldsymbol{U})$. We proceed then to define the closed ball in $\boldsymbol{\mathrm{H}}$
$\mathbf{M}_{R_{1}}=\{\boldsymbol{w}\in\boldsymbol{\mathrm{X}}^{g}:\|\boldsymbol{w}\|_{1,\Omega}\leq R_{1}\},$
and assume that $R_{1}<1$, which (owing to Lemma 3.5 and (3.2)) amounts to consider the assumption on the model data
$\max\{R_{0}^{2}+\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}^{2}+g_{2}+2\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}},$ (32)
$\qquad\frac{1}{\underline{\alpha}}\left[C_{i}\rho_{0}(R_{0}^{2}+\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}^{2})+\mu_{0}\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}+\frac{\underline{\alpha}+\|a\|}{\beta}g_{2}\right](1+\|\theta_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}})+\|\theta_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}\}<1.$
Then we have that $T(\mathbf{M}_{R_{1}})\subseteq\mathbf{M}_{R_{1}}$ ($T$ maps the ball above into itself). We can also assert that $T$ is Lipschitz continuous. By definition, it suffices to verify the Lipschitz continuity of both $S$ (actually, we only require the component $S_{I}$) and $\tilde{S}$.
@@
$b(\boldsymbol{u}_{1}-\boldsymbol{u}_{2},(q,\xi))=\langle\xi,g(\zeta_{1}-\zeta_{2})\rangle_{\Gamma_{\mathrm{m}}}\quad\forall(q,\xi)\in Q\times W.$ (33)
Then, for a given $(q,\xi)\in Q\times W$ and with $\xi\neq 0$, we arrive at {linenomath}
@@
$\lesssim b(\boldsymbol{u}_{1}-\boldsymbol{u}_{2},(q,\xi))$
$=\langle\xi,g(\zeta_{1}-\zeta_{2})\rangle_{\Gamma_{\mathrm{m}}}$
$\leq C\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}}\|\zeta_{1}-\zeta_{2}\|_{1,\Omega},$
where we have used the inf-sup condition from Lemma 3.1, the relation (33), and the Cauchy–Schwarz inequality. Then the result follows after dividing by $\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}}$ on both sides of the inequality. The Lipschitz constant $L_{S}$ depends on the slope of the function $g$ and on the inf-sup constant for $B$.
@@
Then, adding and subtracting $\widetilde{c}(\boldsymbol{w}_{2};\theta_{1},\tau)$ and taking $\tau=\theta_{1}-\theta_{2}$ (which is in $Z_{0}$ since both $\theta_{1},\theta_{2}$ are in $Z$), we obtain {linenomath}
@@ 4 Finite element formulation
In this section we propose a divergence-free finite element scheme to approximate problem (11). The divergence-free requirement is required since we have used the flow incompressibility to write (8). In the following we discuss all properties and stability of the method. Let us consider a shape-regular family of partitions of $\Omega$, denoted by ${\mathcal{T}}_{h}$. We assume that the approximations $\Omega_{h}$ of the domain $\Omega$ is partitioned in simplices such that for $n=2$ we have triangles, whereas tetrahedrons are considered if $n=3$. We denote by $\Gamma_{m,h}$ to the approximation of the membrane boundary $\Gamma_{\mathrm{m}}$. Let $h_{K}$ be the diameter of the element $K\in{\mathcal{T}}_{h}$, and let us define $h:=\max\{h_{K}\,:\,K\in{\mathcal{T}}_{h}\}$.

### v2

@@ 3.1 Well-posedness of the Navier–Stokes equations with membrane boundary conditions
where we have employed (23a) and (23b).
In the context of the fixed-point analysis of (17a)-(17b), for a given $\zeta=\zeta_{0}+\Theta\in Z$ we define the linear functional $G_{\zeta}\in(Q\times W)^{\prime}$ as follows
$\langle G_{\zeta},(q,\xi)\rangle:=\langle\xi,g(\zeta)\rangle_{\Gamma_{\mathrm{m}}}\quad\forall(q,\xi)\in Q\times W.$
Similarly, for a given $\boldsymbol{w}_{0}\in\boldsymbol{\mathrm{H}}_{0}$ we define the linear functional $F_{\boldsymbol{w}_{0},\boldsymbol{U}}\in\boldsymbol{\mathrm{H}}_{0}^{\prime}$:
@@
where $\boldsymbol{U}\in\boldsymbol{\mathrm{H}}^{1}(\Omega)$ is the divergence-free lifting defined in (10). Then, there holds (see [12, Lemma 16])
(24) $\|\boldsymbol{U}\|_{1,\Omega}\lesssim\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}.$
@@
For known liftings $\boldsymbol{U},\Theta$ and given $\boldsymbol{w}_{0}\in\boldsymbol{\mathrm{H}}_{0}$ and $\zeta_{0}\in Z_{0}$, there exists a unique $(\boldsymbol{u}_{0},p,\lambda)\in\boldsymbol{\mathrm{H}}_{0}\times Q\times W$ such that
(25a) $a(\boldsymbol{u}_{0},\boldsymbol{v})+b(\boldsymbol{v},(p,\lambda))$ $=F_{\boldsymbol{w}_{0},\boldsymbol{U}}(\boldsymbol{v})\quad\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{0},$
(25b) $b(\boldsymbol{u}_{0},(q,\xi))$ $=G_{\zeta}(q,\xi)\quad\forall(q,\xi)\in Q\times W.$
@@
(26a) $\|\boldsymbol{u}_{0}\|_{1,\Omega}$ $\leq\frac{1}{\underline{\alpha}}\left[\|F_{\boldsymbol{w}_{0},\boldsymbol{U}}\|_{\boldsymbol{\mathrm{H}}_{0}^{\prime}}+\frac{\underline{\alpha}+\|a\|}{\beta}g_{2}\right],$
(26b) $\|(p,\lambda)\|_{Q\times W}$ $\leq\frac{1}{\beta}\left[\left(1+\frac{\|a\|}{\underline{\alpha}}\right)\|F_{\boldsymbol{w}_{0},\boldsymbol{U}}\|_{\boldsymbol{\mathrm{H}}_{0}^{\prime}}+\frac{\|a\|(\underline{\alpha}+\|a\|)}{\underline{\alpha}\beta}g_{2}\right].$
@@
(27) $\|G_{\zeta}\|_{(Q\times W)^{\prime}}=g_{2}.$
Also, after using triangle inequality together with the boundedness from (12a) and (12d), we obtain
$|F_{\boldsymbol{w}_{0},\boldsymbol{U}}(\boldsymbol{v})|$ $=|-\tilde{a}(\boldsymbol{w}_{0}+\boldsymbol{U},\boldsymbol{w}_{0}+\boldsymbol{U},\boldsymbol{v})-a(\boldsymbol{U},\boldsymbol{v})|$
$\leq C_{i}\rho_{0}\|\boldsymbol{w}_{0}+\boldsymbol{U}\|_{1,\Omega}^{2}\|\boldsymbol{v}\|_{1}+\mu_{0}\|\boldsymbol{U}\|_{1,\Omega}\|\boldsymbol{v}\|_{1,\Omega}$
@@
Let us remark that the unique Stokes velocity from Lemma 3.3 is in the space $\boldsymbol{\mathrm{X}}^{g}$. We now derive a fixed-point strategy for (17a)–(17b). Let us define the map
@@
(28) $0<4R_{0}<1-\sqrt{1-4(\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}^{2}+g_{2}+2\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}})}.$
@@
(29a) $a(\boldsymbol{u}_{01}-\boldsymbol{u}_{02},\boldsymbol{v})+b(\boldsymbol{v},(p_{1}-p_{2},\lambda_{1}-\lambda_{2}))$ $=F_{\boldsymbol{w}_{1},\boldsymbol{U}}(\boldsymbol{v})-F_{\boldsymbol{w}_{2},\boldsymbol{U}}(\boldsymbol{v})\quad\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{0},$
(29b) $b(\boldsymbol{u}_{01}-\boldsymbol{u}_{02},(q,\xi))$ $=0\quad\forall(q,\xi)\in Q\times W.$
@@
On the other hand, in (29) we can use as test functions $\boldsymbol{v}=\boldsymbol{u}_{01}-\boldsymbol{u}_{02}$, $q=p_{1}-p_{2}$, $\xi=\lambda_{1}-\lambda_{2}$ and subtract the two equations to obtain
@@
(30) $L_{\mathcal{J}_{1}}=\frac{1}{\underline{\alpha}^{2}}(2R_{0}+\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}).$
@@
(31) $L_{\mathcal{J}_{1}}=2R_{0}+\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}<\underline{\alpha}^{2}.$
@@ 3.2 Well-posedness of the advection–diffusion equation
(32) $\leq\frac{1}{\underline{\alpha}}\left[\|F_{\boldsymbol{w}_{0},\boldsymbol{U}}\|_{\boldsymbol{\mathrm{H}}_{0}^{\prime}}+\frac{\underline{\alpha}+\|a\|}{\beta}g_{2}\right](1+\|\theta_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}})+\|\theta_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}},$
@@ 3.3 Fixed-point analysis for the coupled flow–transport problem
(33) $\text{find $\boldsymbol{u}=\boldsymbol{u}_{0}+\boldsymbol{U}\in\boldsymbol{\mathrm{H}}$ such that}\ \boldsymbol{u}=T(\boldsymbol{u}),$
where $T:\boldsymbol{\mathrm{H}}\to\boldsymbol{\mathrm{H}}$ is defined as $\boldsymbol{u}\mapsto T(\boldsymbol{u})=({S}_{I}\circ\tilde{S})(\boldsymbol{u}_{0}+\boldsymbol{U})$.
We proceed then to define the closed ball in $\boldsymbol{\mathrm{H}}$ $\mathbf{M}_{R_{1}}=\{\boldsymbol{w}\in\boldsymbol{\mathrm{X}}^{g}:\|\boldsymbol{w}\|_{1,\Omega}\leq R_{1}\}$, and assume that $R_{1}<1$, which (owing to Lemma 3.5 and (3.2)) amounts to consider the assumption on the model data
(34) $\max\left\{R_{0}^{2}+\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}^{2}+g_{2}+2\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}},\right$.
$\qquad\frac{1}{\underline{\alpha}}\left[C_{i}\rho_{0}(R_{0}^{2}+\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}^{2})\left.+\mu_{0}\|\boldsymbol{u}_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}+\frac{\underline{\alpha}+\|a\|}{\beta}g_{2}\right](1+\|\theta_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}})+\|\theta_{\mathrm{in}}\|_{1/2,\Gamma_{\mathrm{in}}}\right\}<1.$
Then we have that $T(\mathbf{M}_{R_{1}})\subseteq\mathbf{M}_{R_{1}}$ ($T$ maps the ball above into itself).
We can also assert that $T$ is Lipschitz continuous. By definition, it suffices to verify the Lipschitz continuity of both $S$ (actually, we only require the component $S_{I}$) and $\tilde{S}$.
@@
(35) $b(\boldsymbol{u}_{1}-\boldsymbol{u}_{2},(q,\xi))=\langle\xi,g(\zeta_{1}-\zeta_{2})\rangle_{\Gamma_{\mathrm{m}}}\quad\forall(q,\xi)\in Q\times W.$
Then, for a given $(q,\xi)\in Q\times W$ and with $\xi\neq 0$, we arrive at
@@
$\lesssim b(\boldsymbol{u}_{1}-\boldsymbol{u}_{2},(q,\xi))=\langle\xi,g(\zeta_{1}-\zeta_{2})\rangle_{\Gamma_{\mathrm{m}}}$
$\leq L_{S}\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}}\|\zeta_{1}-\zeta_{2}\|_{1,\Omega},$
where we have used the inf-sup condition from Lemma 3.1, the relation (35), and the Cauchy–Schwarz inequality. Then the result follows after dividing by $\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}}$ on both sides of the inequality. The Lipschitz constant $L_{S}$ depends on the slope of the function $g$ and on the inf-sup constant for $b(\cdot,\cdot)$.
@@
Then, adding and subtracting $\widetilde{c}(\boldsymbol{w}_{2};\theta_{1},\tau)$ and taking $\tau=\theta_{1}-\theta_{2}$ (which is in $Z_{0}$ since both $\theta_{1},\theta_{2}$ are in $Z$), we obtain
@@ 4 Finite element formulation
In this section we propose a divergence-free FEM to approximate problem (11). This character is required since we have used the flow incompressibility to write (8). In the following we discuss all properties and stability of the method. Let us consider a shape-regular family of partitions of $\Omega$, denoted by ${\mathcal{T}}_{h}$. While the continuous analysis has been conducted for the 2D case, we stress that the construction of the numerical method also holds in 3D. We assume that the approximations $\Omega_{h}$ of the domain $\Omega$ is partitioned in simplices such that for $n=2$ we have triangles, whereas tetrahedrons are considered if $n=3$. We denote by $\Gamma_{m,h}$ to the approximation of the membrane boundary $\Gamma_{\mathrm{m}}$. Let $h_{K}$ be the diameter of the element $K\in{\mathcal{T}}_{h}$, and let us define $h:=\max\{h_{K}\,:\,K\in{\mathcal{T}}_{h}\}$. For the sake of uniformity, in the following we shall use the same notation to denote the FE spaces irrespective of the specific scheme.
@@ 4.1 Divergence-conforming approximation
where the operator $\otimes$ denotes the vector product tensor $\boldsymbol{u}\otimes\boldsymbol{n}]_{ij}=u_{i}\,n_{j},\;1\leq i,j\leq n$. If $e\in\mathcal{E}_{B}$, then we set $\llbracket\boldsymbol{v}\otimes\boldsymbol{n}\rrbracket=\boldsymbol{v}\otimes\boldsymbol{n}$ and $\{\!\!\{\boldsymbol{\tau}\}\!\!\}=\boldsymbol{\tau}$, where $\boldsymbol{n}$ is the unit outward normal vector to $\partial\Omega$.
In the following we specify families of conforming and nonconforming schemes.
4.1.1 Divergence-conforming spaces
Given $k\geq 0$, we define the FE spaces $\boldsymbol{\mathrm{H}}_{h}$, $Q_{h}$, $W_{h}$ and $Z_{h}$ for the velocity, pressure, Lagrange multiplier, and concentration, respectively, by
$\boldsymbol{\mathrm{H}}_{h}$ $=\left\{\boldsymbol{v}_{h}\in\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega)\;:\;\boldsymbol{v}_{h}|_{K}\in[\mathbb{P}_{k+1}(K)]^{2},\,K\in{\mathcal{T}}_{h},\ (\boldsymbol{v}_{h}\cdot\boldsymbol{n})|_{e\in\mathcal{E}_{\mathrm{in}}}=\widehat{u},\ (\boldsymbol{v}_{h}\cdot\boldsymbol{n})|_{e\in\mathcal{E}_{\mathrm{wall}}}=0\right\},$
$\boldsymbol{\mathrm{H}}_{h,0}$ $=\left\{\boldsymbol{v}_{h}\in\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega)\;:\;\boldsymbol{v}_{h}|_{K}\in[\mathbb{P}_{k+1}(K)]^{2},\,K\in{\mathcal{T}}_{h},{\ (\boldsymbol{v}_{h}\cdot\boldsymbol{n})|_{e\in\mathcal{E}_{\mathrm{D}}}=0}\right\},$
$Q_{h}$ $=\left\{q_{h}\in L^{2}(\Omega)\;:\;q_{h}|_{K}\in\mathbb{P}_{k}(K),\,K\in{\mathcal{T}}_{h}\right\},$
$W_{h}$ $=\left\{\xi_{h}\in{\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})}\;:\;\xi_{h}|_{\overline{\Gamma_{j}}}\in\mathbb{P}_{k}(\overline{\Gamma_{j}}),\,j=1,\ldots,n_{\mathcal{E}_{m}}\right\},$
$Z_{h}$ $=\left\{\tau_{h}\in Z\cap C(\overline{\Omega})\;:\;\tau_{h}|_{K}\in\mathbb{P}_{k+1}(K),\,K\in{\mathcal{T}}_{h}\right\}.$
Here, $\mathbb{P}_{r}(\mathcal{O})$, for $r\geq 0$, denotes the space of piecewise polynomials of degree less than or equal to $r$ defined on the entity $\mathcal{O}$, and $\widehat{u}\in\mathbb{P}_{r+1}(\Gamma_{\mathrm{in}})$ is an interpolation of $\boldsymbol{u}_{\mathrm{in}}\cdot\boldsymbol{n}$. For the discrete space of the Lagrange multiplier, we consider a triangulation of $\Gamma_{\mathrm{m}}$ given by $\{\Gamma_{j}\}_{j=1}^{n_{\mathcal{E}_{m}}}$, where $n_{\mathcal{E}_{m}}$ corresponds to the number of facets in $\Gamma_{\mathrm{m}}$. The discrete velocity space is nonconforming in $\boldsymbol{\mathrm{H}}$, and correspond to the well-known divergence-conforming BDM elements family (denoted by $\mathbb{BDM}_{k+1}$) (see [6]). We end this section defining a lowest-order divergence-free $\boldsymbol{\mathrm{H}}^{1}(\Omega)$-nonconforming scheme. For $k=1$ we consider
$\mathbf{CR}_{h}$ $=\{\boldsymbol{v}\in\mathbf{L}^{2}(\Omega)\;:\;\boldsymbol{v}|_{K}\in[\mathbb{P}_{1}(K)]^{2},K\in{\mathcal{T}}_{h},\int_{e}\llbracket\boldsymbol{v}\rrbracket=0,\forall e\in\mathcal{E}_{I}\},$
$\boldsymbol{\mathrm{H}}_{h}$ $=\left\{\boldsymbol{v}_{h}\in\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega)\;:\;\boldsymbol{v}\in\mathbf{CR}_{h},{\ (\boldsymbol{v}_{h}\cdot\boldsymbol{n})|_{e\in\mathcal{E}_{\mathrm{in}}}=\widehat{u},\ (\boldsymbol{v}_{h}\cdot\boldsymbol{n})|_{e\in\mathcal{E}_{\mathrm{wall}}}=0}\right\},$
$\boldsymbol{\mathrm{H}}_{h,0}$ $=\left\{\boldsymbol{v}_{h}\in\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega)\;:\;\boldsymbol{v}\in\mathbf{CR}_{h},{\ (\boldsymbol{v}_{h}\cdot\boldsymbol{n})|_{e\in\mathcal{E}_{\mathrm{D}}}=0}\right\},$
$Q_{h}$ $=\left\{q_{h}\in L^{2}(\Omega)\;:\;q|_{K}\in\mathbb{P}_{0}(K),\,K\in{\mathcal{T}}_{h}\right\},$
$W_{h}$ $=\left\{\xi_{h}\in{\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})}\;:\;\xi|_{\overline{\Gamma_{j}}}\in\mathbb{P}_{0}(\overline{\Gamma_{j}}),\,j=1,\ldots,n_{\mathcal{E}_{m}}\right\},$
$Z_{h}$ $=\left\{\tau_{h}\in Z\cap C(\overline{\Omega})\;:\;\tau|_{K}\in\mathbb{P}_{1}(K),\,K\in{\mathcal{T}}_{h}\right\}.$
Here, the space $\mathbf{CR}_{h}$ is nonconforming in $\boldsymbol{\mathrm{H}}^{1}(\Omega)$ and it corresponds to the lowest-order CR elements (denoted by $\mathbb{CR}_{1}$) which is coupled with piecewise constants for the pressure (see [13, 17]).
4.1.2 The discrete div-conforming problem
As the discrete velocity now lives in $\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega)$ and its normal trace is in $\mathrm{H}^{-1/2}(\partial\Omega)$, the pairings on $\Gamma_{\mathrm{m}}$ from (7),(9) suggest a discrete Lagrange multiplier space conforming with $\mathrm{H}^{1/2}(\Gamma_{\mathrm{m}})$ instead of $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$. In that case, in (9) one should use $\mathcal{R}^{-1}_{1/2}(g(\theta))$ instead of $g(\theta)$ (where $\mathcal{R}_{1/2}$ denotes the Riesz map between $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$ and its dual). However, we maintain $W_{h}$ defined as conforming with $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$ as in the previous section. We bear in mind that the off-diagonal bilinear form (to be denoted $b_{h}(\cdot,\cdot)$ in (37) below) is therefore slightly different, needing the Riesz representative of the Lagrange multiplier.